# IDESplat: Iterative Depth Probability Estimation for Generalizable 3D Gaussian Splatting

### v2

@@ 3.3 Gaussian Focused Module
For the remaining Gaussian parameters, we introduce a window-based Gaussian Focused Module that can filter irrelevant Gaussians and retains the most relevant tokens for attention weight computation. The window-based attention performs pairwise interactions for each Gaussian token, often including numerous irrelevant ones. This dense interaction not only slows down the model but also introduces noise into the attention results. Inspired by [24], we introduce a Gaussian Focused Layer that reuses the previous layer’s Gaussian correlation map to guide attention in the current layer. Formally, for the Gaussian parameters of a given view, $G\in\mathbb{R}^{C\times H\times W}$, we apply three linear layers to obtain $\mathbf{Q}$, $\mathbf{K}$, and $\mathbf{V}$. We use a matrix $\mathbf{I}_{G}$ to record the indices of Gaussian tokens with high similarity. $\mathbf{I}_{G}^{0}$ is initialized as an all-one matrix and is updated after each similarity computation. Then, we compute the Gaussian similarity map as follows:
@@
3.4 Model Architecture
Our IDESplat consists of three components: the feature extraction backbone, the iterative depth estimation process, and the gaussian focused module.
Feature Extraction Backbone. The feature extraction backbone follows the design of  [46] and can be divided into two parts: multi-view feature extraction and monocular feature extraction. The pre-trained Unimatch [43] network is used as the multi-view feature extraction branch, and the pre-trained DepthAnything V2 [49] network based on the ViT-small version is used as the monocular feature extraction branch. The outputs of these two branches are directly stacked along the channel dimension using concatenation and then fused through convolution. This backbone captures rich multi-view geometry and texture information, which is then used as input to the iterative depth estimation process.
Iterative Depth Estimation Process. To achieve both high-quality reconstruction performance and real-time efficiency, we design an iterative pipeline with three Depth Probability Boosting Units (DPBUs). As a result, the pipeline produces three progressively refined depth outputs. Each DPBU consists of two Warp-Index Epipolar Attention layers connected in sequence, and their attention maps are combined via a Hadamard product to multiplicatively boost depth probabilities and yield accurate depth estimates. Across the entire pipeline, IDESplat performs six warp operations. In addition, the feature resolution increases stage by stage—the input resolutions of the three units are $64\times 64$, $128\times 128$, and $256\times 256$, respectively.
Gaussian Focused Module. The Gaussian Focused Module consists of six Gaussian-focused layers, with a feature partition window size of 16 and a shift window strategy. After each attention calculation, sparse operations are applied to retain only the most relevant half of the Gaussian positions. The number of retained Gaussian weight values in each layer is [256, 256, 128, 128, 64, 64], respectively. The module uses multi-head attention with 6 attention heads, and the total number of channels is 256. Additionally, we incorporate LePE positional encoding [7] during the attention computation.
@@ 4.1 Experiment Setting
Implementation details. For a fair comparison, we followed the commonly used training setup [5, 22, 46]. Specifically, our training experiments were conducted on 8 RTX 4090 GPUs with a total batch size of 16, using the AdamW [25] optimizer for 300,000 iterations. We employed a cosine learning rate schedule to optimize our model. For the pre-trained Depth Anything V2 [49] backbone, we used a learning rate of $2\times 10^{-6}$, while the remaining layers were trained with a learning rate of $2\times 10^{-4}$.
@@ 4.2 Main Results
IDESplat (Ours) 37.6 27.80 0.893 0.108 29.04 0.866 0.124
Quantitative results. As shown in Tab. 1, our IDESplat achieves state-of-the-art performance on all visual quality metrics in both the RealEstate10K and ACID benchmarks. Specifically, on the RE10K dataset, compared to DepthSplat, our method improves PSNR by 0.33 dB, while using only 10.7% of its parameters. Additionally, compared to MonoSplat, which has a similar number of parameters, our method achieves a significant 1.12 dB improvement. For SSIM and LPIPS metrics, our method also achieves the best results of 0.893 and 0.108, respectively. On the ACID dataset, our method achieves a 0.41 dB improvement in PSNR over MonoSplat, reaching a maximum of 29.04 dB. The superior performance of IDESplat can be attributed to its iterative boosting design for depth probability prediction, which gradually refines the depth map. This design allows our method to achieve reliable and stable depth estimation results with fewer parameters, enabling accurate Gaussian mean parameter prediction and high-quality scene reconstruction.
@@
Visual comparison of depth maps. We conducted a direct qualitative comparison of depth maps predicted by IDESplat and existing state-of-the-art models, including MVSplat [5], MonoSplat [22], and DepthSplat [46]. It is important to note that the generalizable 3DGS task typically does not have depth ground truth, so we analyzed the results using the corresponding reference RGB images as input. As shown in Fig. 4, our proposed IDESplat method consistently outperforms existing methods in both indoor and outdoor scenes. Even in areas where the foreground and background are highly similar or in regions with rich texture details, our method can clearly distinguish objects at different depths. Additionally, the depth estimation results of IDESplat exhibit better and more realistic depth effects, with finer texture details, in both indoor and outdoor scenes. This is because our method can integrate the results of feature similarity from multiple warps using the depth probability boosting strategy, enabling more accurate depth estimation. Furthermore, by performing warp calculations at the original scale, our method captures and utilizes fine textures in the image, leading to more detailed and precise predictions. More visual results can be found in the supplementary materials.
@@ 4.3 Ablation Study
To comprehensively analyze the efficiency and reconstruction performance of IDESplat with different iteration counts, we conducted detailed ablation experiments. In these experiments, the Gaussian Focused Module and Depth Probability Boosting Strategy were used by default. Additionally, during the iterative process, we gradually increased the resolution, and the depth search range was progressively halved. For 3 iterations, the feature resolutions during warping were $\frac{1}{4}$, $\frac{1}{2}$, and $1$ of the original resolution. For 4 iterations, the model iterates twice at the largest feature size. When the number of iterations is 0, it represents performing a single warp for depth probability estimation. The results of the experiment, as shown in Tab. 5, indicate that even with a single iteration, our method provides a 0.45 dB improvement. With 3 iterations, the performance improves by 0.93 dB. Furthermore, as the number of iterations increases, the increase in parameters is minimal, and the memory usage remains comparable to existing methods. Although inference time increases with more iterations, the performance improvement is substantial by the third iteration, while still maintaining real-time inference.
@@ A. Training Details
For a fair comparison, we trained our IDESplat using a standard setup [5, 22, 46]. The training was done on 8 RTX 4090 GPUs with a batch size of 16, using the AdamW [25] optimizer for 300,000 iterations, taking almost 3 days. We used a cosine learning rate schedule. For the pre-trained Depth Anything V2 [49] backbone, the learning rate was $2\times 10^{-6}$, while other layers were trained with a learning rate of $2\times 10^{-4}$. The network was trained with a combination of MSE and LPIPS losses between the rendered and ground truth images. Following [46], for the newly added DL3DV dataset, we trained at a resolution of $256\times 448$. First, we pre-trained on re10k and then fine-tuned on the DL3DV dataset for 100K iterations, with a total batch size of 4, and the number of input views was randomly sampled from 2 to 6. During inference, we evaluated the model’s performance on different numbers of input views.
@@ C. More Experimental Results
Table 7: Ablation results for DPBU with different numbers of Warp-Index Epipolor Attention. All results are reported on the RealEstate10K dataset.

### v3

@@ 3.3 Gaussian Focused Module
For the remaining Gaussian parameters, we introduce a window-based Gaussian Focused Module that can filter irrelevant Gaussians and retains the most relevant tokens for attention weight computation. The window-based attention performs pairwise interactions for each Gaussian token, often including numerous irrelevant ones. This dense interaction not only slows down the model but also introduces noise into the attention results. Inspired by [23], we introduce a Gaussian Focused Layer that reuses the previous layer’s Gaussian correlation map to guide attention in the current layer. Formally, for the Gaussian parameters of a given view, $G\in\mathbb{R}^{C\times H\times W}$, we apply three linear layers to obtain $\mathbf{Q}$, $\mathbf{K}$, and $\mathbf{V}$. We use a matrix $\mathbf{I}_{G}$ to record the indices of Gaussian tokens with high similarity. $\mathbf{I}_{G}^{0}$ is initialized as an all-ones matrix and is updated after each similarity computation. Then, we compute the Gaussian similarity map as follows:
@@ 4.1 Experiment Setting
Implementation details. For a fair comparison, we followed the commonly used training setup [4, 21, 46]. Specifically, our training experiments were conducted on 8 RTX 4090 GPUs with a total batch size of 16, using the AdamW [24] optimizer for 300,000 iterations. We employed a cosine learning rate schedule to optimize our model. For the pre-trained Depth Anything V2 [50] backbone, we used a learning rate of $2\times 10^{-6}$, while the remaining layers were trained with a learning rate of $2\times 10^{-4}$. Following DepthSplat [46], we train our model with MSE and LPIPS losses.
@@ 4.2 Main Results
IDESplat (Ours) 37.6 27.80 0.893 0.108 28.94 0.866 0.130
Quantitative results. As shown in Tab. 1, our IDESplat achieves state-of-the-art performance on all visual quality metrics in both the RealEstate10K and ACID benchmarks. Specifically, on the RE10K dataset, compared to DepthSplat, our method improves PSNR by 0.33 dB, while using only 10.7% of its parameters. Additionally, compared to MonoSplat, which has a similar number of parameters, our method achieves a significant 1.12 dB improvement. For SSIM and LPIPS metrics, our method also achieves the best results of 0.893 and 0.108, respectively. On the ACID dataset, our method achieves a 0.31 dB improvement in PSNR over MonoSplat, reaching a maximum of 28.94 dB. The superior performance of IDESplat can be attributed to its iterative boosting design for depth probability prediction, which gradually refines the depth map. This design allows our method to achieve reliable and stable depth estimation results with fewer parameters, enabling accurate Gaussian mean parameter prediction and high-quality scene reconstruction.
@@
Visual comparison of depth maps. We conducted a direct qualitative comparison of depth maps predicted by IDESplat and existing state-of-the-art models, including MVSplat [4], MonoSplat [21], and DepthSplat [46]. It is important to note that the generalizable 3DGS task typically does not have depth ground truth, so we analyzed the results using the corresponding reference RGB images as input. As shown in Fig. 5, IDESplat consistently outperforms existing methods in both indoor and outdoor scenes. Even in regions with similar foreground and background appearances or rich texture details, it can better distinguish objects at different depths and produce more realistic depth estimates with finer details. This advantage comes from aggregating feature similarity across multiple warps via the depth probability boosting strategy, as well as performing warp computation at the original scale to better preserve fine textures. More visual results are provided in the supplementary materials.
@@
Figure 4: Qualitative comparison of depth error maps on the ScanNet dataset.
Depth estimation experiment on ScanNet. We report depth estimation results on the ScanNet dataset in Table 4. IDESplat consistently outperforms UniMatch and DepthSplat across standard quantitative depth estimation metrics. For qualitative comparison, we further visualize the depth error maps on the ScanNet dataset in Fig. 4. Compared with the baseline methods, IDESplat exhibits fewer large-error regions and produces more accurate depth estimates, especially around geometrically complex areas.
Table 4: Depth estimation results on the ScanNet dataset.
Method Abs Rel $\downarrow$ RMSE $\downarrow$ RMSElog $\downarrow$
UniMatch 0.059 0.179 0.082
DepthSplat 0.045 0.125 0.061
IDESplat 0.039 0.116 0.053
@@ 4.3 Ablation Study
We conduct ablation experiments to evaluate the efficiency and reconstruction performance of IDESplat under different iteration counts. In these experiments, the Gaussian Focused Module and Depth Probability Boosting Strategy were used by default. Additionally, during the iterative process, we gradually increased the resolution, and the depth search range was progressively halved. For 3 iterations, the feature resolutions during warping were $\frac{1}{4}$, $\frac{1}{2}$, and $1$ of the original resolution. For 4 iterations, the model iterates twice at the largest feature size. When the number of iterations is 0, it represents performing a single warp for depth probability estimation. The results of the experiment, as shown in Tab. 6, indicate that even with a single iteration, our method provides a 0.45 dB improvement. With 3 iterations, the performance improves by 0.93 dB. Increasing the iteration count adds little parameter overhead and keeps memory usage comparable to existing methods. Despite the longer inference time, 3 iterations already provide a substantial gain while maintaining real-time inference.
@@ 5 Conclusion
Acknowledgement
This work was supported by the National Natural Science Foundation of China under Grant No. 62476051.
@@ A. Training Details
For a fair comparison, we trained our IDESplat using a standard setup [4, 21, 46]. The training was done on 8 RTX 4090 GPUs with a batch size of 16, using the AdamW [24] optimizer for 300,000 iterations, which took approximately 3 days. We used a cosine learning rate schedule. For the pre-trained Depth Anything V2 [50] backbone, the learning rate was $2\times 10^{-6}$, while other layers were trained with a learning rate of $2\times 10^{-4}$. The network was trained with a combination of MSE and LPIPS losses between the rendered and ground truth images. Following [46], for the newly added DL3DV dataset, we trained at a resolution of $256\times 448$. First, we pre-trained on RE10k and then fine-tuned on the DL3DV dataset for 100K iterations, with a total batch size of 4, and the number of input views was randomly sampled from 2 to 6. During inference, we evaluated the model’s performance on different numbers of input views.
@@ C. More Experimental Results
Table 8: Ablation results for DPBU with different numbers of Warp-Index Epipolar Attention. All results are reported on the RealEstate10K dataset.